OpenStudy (anonymous):

How do I evaluate ^3Sqrt-64

OpenStudy (tkhunny):

Guess. (-2)^3 = (-2)(-2)(-2) = -8 -- Too low. (-3)^3 = (-3)(-3)(-3) = -27 -- Too low. (-4)^3 = (-4)(-4)(-4) = -64 -- Tripped over it!